OpenStudy (anonymous):

hi will you guys help me with this please this is the exercises and i can figured out what to do 5/10 = / 4/16

OpenStudy (anonymous):

i think it has something to do with having the same common denometer?

OpenStudy (anonymous):

did that help you?

OpenStudy (anonymous):

alright chris first you are looking for a fraction... let's name this fraction X

OpenStudy (anonymous):

whenever you are trying to find an answer name it a variable this will help you solve the problem

OpenStudy (anonymous):

so you'll get \[\frac{5}{10}=\frac{x}{\frac{4}{16}}\]

OpenStudy (anonymous):

so far do you understand?

OpenStudy (anonymous):

chris?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright so do you know how to simplify the right side?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright so what would you do

OpenStudy (anonymous):

to get rid of the fraction in the denominator

OpenStudy (anonymous):

so the simplify fraction is 1/2

OpenStudy (anonymous):

\[\frac{x}{\frac{4}{16}}=\frac{x}{1}*\frac{16}{4}\]

OpenStudy (anonymous):

do you know this property?

OpenStudy (anonymous):

i dont get it for real not really this was never explained to me really thank you dude

OpenStudy (anonymous):

the division of a numerator and a fraction denominator is the same as the multiplication of it's reciprocal denominator

OpenStudy (anonymous):

so say you have something like\[\frac{\frac{a}{b}}{\frac{c}{d}}=\frac{a}{b}*\frac{d}{c}\]

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

yes i do get that but in this case i have to find the fraction that divided by 4/16 will be equal to 5/10 is not lkike find the answer i do know how to divided a fraction but this is kind of tricki

OpenStudy (anonymous):

in this case x is not a fraction but you can make it a fraction by putting it over a one \[\frac{x}{1}=x\]

OpenStudy (anonymous):

well we haven't finishd yet haha i gurantee that we will find the answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{5}{10}=\frac{16x}{4}\]

OpenStudy (anonymous):

all i did was multiply the right side out . See?

OpenStudy (anonymous):

you multipy by 1

OpenStudy (anonymous):

So i used that rule to get rid of the fraction in the denominator \[\frac{5}{10}=\frac{\frac{x}{1}}{\frac{4}{16}}=\frac{x}{1}*\frac{16}{4}=\frac{16x}{4}\]

OpenStudy (anonymous):

alright?

OpenStudy (anonymous):

or do you still not understand how that property works?

OpenStudy (anonymous):

and yes it is known as multiplying by one

OpenStudy (anonymous):

yes but in the exercises my teacher wrote like this 5/10= / 4/16

OpenStudy (anonymous):